Show that if π‘Žπ‘‘βˆ’π‘π‘ β‰  0, then the function 𝑓(π‘₯)=π‘Žπ‘₯+𝑏/𝑐π‘₯+𝑑 is one-to-one and find its inverse.


1
Expert's answer
2021-10-19T15:44:01-0400
"f(x)=\dfrac{ax+b}{cx+d}"

"cx+d\not=0=>x\not=-d/c"

Domain: "(-\infin, -d/c)\cup(-d/c, \infin)"

Replace "f(x)" with "y"


"y=\dfrac{ax+b}{cx+d}"

Switch "x" and "y"


"x=\dfrac{ay+b}{cy+d}"

Solve for "y"


"cxy+dx=ay+b"

"y=\dfrac{b-dx}{cx-a}"

"cx-a\not=0=>x\not=c/a"


"f^{-1}\circ f=\dfrac{b-d(\dfrac{ax+b}{cx+d})}{c(\dfrac{ax+b}{cx+d})-a}"

"=\dfrac{bcx+bd-adx-bd}{acx+bc-acx-ad}"

"=\dfrac{bcx-adx}{bc-ad}=x, bc\not=ad=>ad-bc\not=0"

Therefore Β if "ad-bc\not=0," then the function "f(x)=\dfrac{ax+b}{cx+d}" is one-to-one and find its inverse.